if 2x+3=4.find x


1
Expert's answer
2021-06-14T15:43:40-0400
"2x+3=4"

"2x+3-3=4-3"


"2x=1"

"\dfrac{2x}{2}=\dfrac{1}{2}"

"x=\dfrac{1}{2}"